What is the domain? p(x)=x^3-x^2-6. Is it the set {x|x is a real number and x does not equal 6}?
Answer by nabla(475)

Polynomials have all real numbers as their domains. The question of "what is the domain" is really one asking where is the function defined.
Put 6 into the function. Do you get a "normal" function value? Yes, p(6)=174.

For another example, consider f(x)=1/x. Here we know that the function behaves "funny" around x=0. Moreover, many people are aware of the fact that "we can not divide by zero." Although it takes a bit more math to see why this is true concisely, take a look at the graph and you will see some strange behavior.

A final example, consider g(x)=ln(x). This is the natural log function. ln(x)=y means the same thing as e^y=x. Hence, x is always positive and not 0 as e^y=0 is nonsense.